OpenStudy (anonymous):

Which of the following is the simplified form of fifth root of x times the fifth root of x times the fifth root of x times the fifth root of x?

OpenStudy (***[isuru]***):

u mean \[\large\ 5\sqrt{x} \times 5\sqrt{x} \times 5\sqrt{x} \times 5\sqrt{x} \] ?

OpenStudy (anonymous):

Yes! @***[ISURU]***

OpenStudy (***[isuru]***):

so... \[\large\ a \times a \times a \times a = a^4\] if u think \[\large\ a = 5 \sqrt{x}\] then its \[\large\ (5\sqrt{x})^4 = 5^4 \times (\sqrt{x})^4 = 5^4 \times (x^\frac{ 1 }{ 2 })^4 = 5^4 \times x^{4 \times \frac{ 1 }{ 2 }} \]\[\large\ = 5^4 \times x^2 \]

OpenStudy (***[isuru]***):

got it ?

OpenStudy (anonymous):

this are the only answer choices: a) x to the 1 over fifth power b) x to the 4 over fifth power c) x to the 4 over twentieth power d) x

OpenStudy (anonymous):

is it b?

OpenStudy (***[isuru]***):

wait ...

OpenStudy (***[isuru]***):

u mean a question like this.. rnt u ? \[\large\ \sqrt[5]{x} \times \sqrt[5]{x} \times \sqrt[5]{x} \times \sqrt[5]{x}\]

OpenStudy (***[isuru]***):

these r 2 different notations \[\large\ 5 \sqrt{x} = 5 \times \sqrt{x}\] but \[\large\ \sqrt[5]{x} = x^\frac{ 1 }{ 5 } \]

OpenStudy (anonymous):

thankyou! , can u help with one more?

OpenStudy (***[isuru]***):

sure, but did u find the right answer for this one ?

OpenStudy (anonymous):

yeah. the other one is Which of the following is the radical expression of 4 times d to the three eighths power? A) 4 times the eighth root of d to the third power B) 4 times the third root of d to the eighth power C) eighth root of 4 d to the third power D) third root of 4 d to the eighth power

OpenStudy (***[isuru]***):

so this one speaks about... \[\large\ 4 \times (d)^\frac{ 3 }{ 8 }\] ur first choice says\[\large\ 4 \times \sqrt[8]{d} \] and the second is \[\large\ 4 \times (\sqrt[3]{d})^8\] the third one is\[\large\ (\sqrt[8]{4d})^3\] and the last one is\[\large\ (\sqrt[3]{4d})^8\] now ...what do u think ?

OpenStudy (***[isuru]***):

srry my mistake the first choice is\[\large\ 4 \times (\sqrt[8]{d})^3\]

OpenStudy (***[isuru]***):

heres the help u need\[\large\ d^\frac{ 3 }{ 8 } = d^{\frac{ 1 }{ 8 } \times 3} = (d^\frac{ 1 }{ 8 })^3 = (\sqrt[8]{d})^3\] now u will be able to find out ur answer...and srry for the late reply.. i was busy with some work